A grating spectrometer is used to resolve first order peaks from two wavelengths of ?1=555.00 nm and ?2=555.29 nm. What is the minimum number of slits this grating must have to resolve these wavelengths?
Answer #1

as we know resolving power

R = L / dL

for first order resolving power

1 * N = 555 / ( 555.29 - 555)

N = 1913.8

N ≈ 1914

=========